OpenStudy (anonymous):

Jonathon has a bag full of 14 balls to sell. He sells the baseballs for $2 dollars and the volleyballs for $5 and earns a total of $43. If the baseballs are represented by x and the volleyballs are represented by y, which of the following systems of equations can be used to calculate the number of baseballs and volleyballs sold?

OpenStudy (anonymous):

The first equation is x + y = 14

OpenStudy (anonymous):

baseballs are x, volleyballs are y, 14 is the total number of balls

OpenStudy (anonymous):

the second equation is 2x + 5y = 43

OpenStudy (anonymous):

2x because baseballs are 2 dollars, 5y because volleballs are 5 dollars, and 43 becuase of the total amount

OpenStudy (anonymous):

now to solve for x and y, you need to isolate for either x or y in the first equation

OpenStudy (anonymous):

x+y=14 get x or y by itself